A tree casts a shadow of 12 feet. At the same time a 6-foot man casts a 3-foot shadow. How tall is the tree?

Answer by checkley79(3341)

12/X=3/6 CROSS MULTIPLY.
3X=12*6
3X=72
X=72/3
X=24 FEET TALL.